# AF Embeddability of the C*-Algebra of a Deaconu-Renault Groupoid

### v1

@@ 4.1. Homology in algebra
First we present the definition of homology groups from abstract algebra. A possible reference is [16, Section 17.1].
@@ 4.2. Homology for groupoids
We define homology groups for Haudorff, second countable, ample groupoids, as in [20, 39]. We fix $X,Y$ to be locally compact, Hausdorff, second countable, totally disconnected spaces.
Here $C_{c}(X,\mathbb{Z})$ denotes the set of continuous and compactly supported functions on $X$ assuming integer values. Given $f\in C_{c}(X,\mathbb{Z})$ and a local homeomorphism $\sigma:X\rightarrow Y$, define for every $y\in Y$,
@@
The group $H_{0}(G)$ is not necessarily an ordered group because the intesection $H_{0}(G)^{+}\cap(-H_{0}(G)^{+})$ may be different from $\{0\}$. See [39, Remark 3.2].
@@
We prove the lemma below, stated in [39, Subsection 3.2] for étale groupoids.
@@
Let $\rho:\rightarrow H$ be a homomorphism that is also a local homeomorphism. Then, for each $n\geq 0$, $\rho$ induces a homomorphism $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ defined by
@@
In this paper, for a function $\tau:C_{c}(G^{(0)},\mathbb{Z})\rightarrow C_{c}(H^{(0)},\mathbb{Z})$ we use the notation $[\tau]$ to denote the function $[\tau]:H_{0}(G)\rightarrow H_{0}(H)$ given by $[\tau]([f])=[\tau(f)]$. Moreover, for a homomorphism $\rho:G\rightarrow H$ that is also a local homeomorphism, we have $H_{0}(\rho)=[\rho_{\ast}^{(0)}]$.
@@
[39, Proposition 3.5] Let $n\geq 1$. If $\rho,\sigma:G\rightarrow H$ are similar homomorphisms and local homeomorphisms, then $H_{n}(\rho)=H_{n}(\sigma)$.
@@
[39, Proposition 3.5] If $G$ and $H$ are homologically similar, then they have isomorphic homology groups. In particular, if $n\geq 1$, $\rho:G\rightarrow H$ and $\sigma:H\rightarrow G$ are local homeomorphisms and homomorphisms satisfying (4.2), then $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ is an isomorphism with inverse $H_{n}(\sigma)$. Moreover, both $H_{0}(\rho)$ and $H_{0}(\sigma)$ preserve positive elements.
4.3. Homology for AF groupoids
Like K-theory, homology groups are continuous with respect to inductive limits. We start this subsection by looking at the building blocks of AF groupoids, i.e., groupoids of the form $R(\sigma)$ and finding a property that helps us find the equivalence in $H_{0}(R(\sigma))$.
@@ 4.3. Homology for AF groupoids
5. K-theory for AF Groupoids
Here we extend the isomorphism [20, Corollary 5.2] between the K-theory and the homology group for AF groupoids, showing that the map preserves positive elements, and giving a formula for the isomorphism.
Although it is often difficult to find the K-theory of an arbitrary C*-algebra, finding the K-theory of $M_{n}$ is a easy. For this case, we get the isomorphism $K_{0}(M_{n})\rightarrow\mathbb{Z}$ given by $[p]_{0}\mapsto\mathrm{tr}\hphantom{.}p$, where $\mathrm{tr}\hphantom{.}p$ is the matrix trace of $p\in\mathcal{P}_{\infty}(M_{n})$. Similarly, for a locally compact, Hausdorff, second countable, totally disconnected space $X$, the K-theory of $C_{0}(X,M_{n})$ is also described in terms of a trace operator.
To find the K-theory of an AF groupoid, we first consider the building blocks of the groupoid. In other words, we consider an elementary groupoid of the form $R(\sigma)$. We show that $R(\sigma)$ is locally approximated by groupoids of the form $C_{0}(V,M_{n})$. Then we apply approximation to find a formula for the $K_{0}$-group of $C^{*}(R(\sigma))$ and an isomorphism from this group to the corresponding homology group.
@@ 5. K-theory for AF Groupoids
Let $p\in C_{0}(X,M_{n})$ be a projection.Let $K\subset X_{\varphi}$ be a compact subset such that $\|p(x)\|<1$ for every $x\notin K$. It follows from [51, Propositions 2.2.4 and 2.2.7] that $\mathrm{tr}\hphantom{.}p(x)=0$ for all $x\notin K$. Then $\mathrm{tr}\hphantom{.}p\in C_{c}(X_{\varphi},\mathbb{N})$ and $p\in C_{c}(X,\mathbb{N})$.
@@
Then $\widetilde{p}$ and $\widetilde{q}$ are projections in $C_{0}(X,M_{n})$. Also, $\widetilde{p}=v^{*}\widetilde{q}v$ and $v^{*}v=vv^{*}=\mathrm{id}1_{K}$. Considering the unital C*-algebra $C(K,M_{n})\subset C_{0}(X,M_{n})$, it follows from [51, Propositions 2.2.4 and 2.2.7] that there exists $u_{p},u_{q}\in C_{0}(X,M_{n})$ with support in $K$ such that
@@
    where we apply Renualt’s $j$ map [48, Proposition II.4.2] to identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$.
@@
    By [55, Proposition 9.3.1], for all $x\in X$, the $\ast$-representation $\pi_{x}:C_{c}(R(\sigma))\rightarrow B(\ell^{2}(R(\sigma)_{x}))$ is given by
@@
    Given $x\in X$, let $u=\varphi(\sigma(x))\in X_{\varphi}$. [55, Proposition 9.3.1] implies that there exists a unitary $U_{(x,u)}:\ell^{2}(R(\sigma)_{u})\rightarrow\ell^{2}(R(\sigma)_{x})$ such that $\pi_{x}=U_{(x,u)}\pi_{u}U_{(x,u)}^{*}$. Then $\|\pi_{x}(f)\|=\|\pi_{u}(f)\|$. Therefore,
@@
The following theorem gives an isomorphism of the ordered groups $K_{0}(C^{*}(G))$ and $H_{0}(G)$. Corollary 5.2 by Farsi, Kumjian, Pask and Sims [20] and Matui’s [39, Theorem 4.10] prove isomorphisms of homology groups of groupoids and the corresponding $K_{0}$-group. However, [20, Corollary 5.2] does not prove whether such isomorphism preseves positive elements from one group to another, and Matui only considers groupoids with compact unit spaces, without providing a detailed proof. Our theorem generalises both results, and we describe the isomorphism by an explicit formula.
@@
    the isomorphism $\nu$ is precisely the map given by [20, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
@@ 3. Part 3
6. The Commutative Diagram
@@ 6.1. Part 1 of the diagram
6.2. Part 2 of the diagram
In the second part of the diagram, we have an isomorphism between homology groups of different groupoids. We will obtain this isomorphism by applying the techniques of homological similarity of groupoids. Let us define the groupoid $c^{-1}(0)$.
Until the rest of this section, we fix $\varphi:X\rightarrow X$ be a continuous section of the surjective local homeomorphism $\sigma:X\rightarrow X$ that defines $\mathcal{G}$. Let us define the map that induces the similarity of the groupoids $\mathcal{G}(c)$ and $c^{-1}(0)$.
@@ 6.2. Part 2 of the diagram
7. Main Result
The main result of this paper is Theorem 1, which gives a condition on the map $\sigma:X\rightarrow X$, characterising when the corresponding groupoid C*-algebra $C^{*}(\mathcal{G})$ is AF embeddable. We prove this result by applying the commutative diagram of Section 4 in combination with Theorem 1.1.
@@ 7. Main Result
Let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the continuous cocycle given by $c(x,k,y)=k$ for $(x,k,y)\in\mathcal{G}$. By [20, Corollary 5.2], we have that $\mathcal{G}(c)$ is an AF groupoid, and therefore the C*-algebra $C^{*}(G(c))$ is AF. For simplicity of notation, we denote this C*-algebra by $B$.
@@ 8.1. Graph algebras
Given a row-finite graph with no sources, [33, Theorem 4.2] gives an isomorphism from the graph algebra $C^{*}(E)$ to the C*-algebra $C^{*}(G_{E})$ defined by $1_{e}\mapsto 1_{Z(e,r(e))}$, where
@@
The groupoid is almost identical to the groupoid in [33, Definition 2.3], but we made some changes to make it compatible with the notation in Raeburn’s book [46]. In constrast to our definition, in [33] two consecutive edges $x_{i},x_{i+1}$ satisfy $r(x_{i})=s(x_{i+1})$. Also, their groupoid has elements of the form $(x,-k,y)$ instead of $(x,k,y)$.
@@
    But if $y\neq z_{k+1}$ and $\sigma^{n}(y)=z_{k}$, then $y$ satisfies condition $(2)$ of Lemma 8.1 for $n,p\geq 1$. So $f(y)\leq 0$ by item (b) of this proof. Hence, $f(z_{k})\leq f(z_{k+1})$. It is straightforward from the definition that $z_{p}=z_{0}$. This implies that $f(z_{0})=f(z_{1})=\dots=f(z_{p})$. By (8.1), $f(y)=0$ for all $y\neq z_{1}$ such that $\sigma^{n}(y)=z_{0}=\sigma^{n}(x)$. Since $z_{1}$ satisfies condition $(2)$ of 8.1 for 00 and $p$, we have that $x\neq z_{1}$. Therefore $f(x)=0$.
@@
So far we have proved $(i)\Leftrightarrow(v)$. By Theorem 1, we have the equivalence $(i)\Leftrightarrow(ii)\Leftrightarrow(iii)\Leftrightarrow(v)$. The implication $(iii)\Rightarrow(iv)$ is straightforward, and $(iv)\Rightarrow(v)$ follows from the same argument used in [54, Theorem 1.1]. ∎

### v2

@@ 2.3. Homology in algebra
First we present the definition of homology groups from abstract algebra. A possible reference is [11, Section 17.1].
@@ 2.4. Homology for groupoids
We define homology groups for Hausdorff, second countable, ample groupoids, as in [12, 20]. We fix $X,Y$ to be locally compact, Hausdorff, second countable, totally disconnected spaces.
Here, $C_{c}(X,\mathbb{Z})$ denotes the set of continuous and compactly supported functions on $X$ assuming integer values. Given $f\in C_{c}(X,\mathbb{Z})$ and a local homeomorphism $\sigma:X\rightarrow Y$, define for every $y\in Y$,
@@
The group $H_{0}(G)$ is not necessarily an ordered group because the intesection $H_{0}(G)^{+}\cap(-H_{0}(G)^{+})$ may be different from $\{0\}$. See [20, Remark 3.2].
@@
We prove the lemma below, stated in [20, Subsection 3.2] for étale groupoids.
@@
Let $\rho:G\rightarrow H$ be a homomorphism that is also a local homeomorphism. Then, for each $n\geq 0$, $\rho$ induces a homomorphism $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ defined by
@@
In this paper, for a function $\tau:C_{c}(G^{(0)},\mathbb{Z})\rightarrow C_{c}(H^{(0)},\mathbb{Z})$ such that $f_{1}\sim f_{2}$ in $H_{0}(G)$ imples $\tau(f_{1})\sim\tau(f_{2})$ in $H_{0}(H)$, we use the notation $[\tau]$ to denote the function $[\tau]:H_{0}(G)\rightarrow H_{0}(H)$ given by $[\tau]([f])=[\tau(f)]$. Moreover, for a homomorphism $\rho:G\rightarrow H$ that is also a local homeomorphism, we have $H_{0}(\rho)=[\rho_{\ast}^{(0)}]$.
@@
[20, Proposition 3.5] Let $n\geq 1$. If $\rho,\sigma:G\rightarrow H$ are similar homomorphisms and local homeomorphisms, then $H_{n}(\rho)=H_{n}(\sigma)$.
@@
[20, Proposition 3.5] If $G$ and $H$ are homologically similar, then they have isomorphic homology groups. In particular, if $n\geq 1$, $\rho:G\rightarrow H$ and $\sigma:H\rightarrow G$ are local homeomorphisms and homomorphisms satisfying (2.4), then $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ is an isomorphism with inverse $H_{n}(\sigma)$. Moreover, both $H_{0}(\rho)$ and $H_{0}(\sigma)$ preserve positive elements.
3. Homology for AF groupoids
In this section, we study the properties of the homology groups of AF groupoids. Like K-theory, homology groups are continuous with respect to inductive limits. We start this subsection by looking at the building blocks of AF groupoids, i.e., groupoids of the form $R(\sigma)$ and finding a property that helps us find the equivalence in $H_{0}(R(\sigma))$.
@@ 3. Homology for AF groupoids
4. Main result
Here we extend the isomorphism [12, Corollary 5.2] between the K-theory and the homology group for AF groupoids, showing that the map preserves positive elements, and giving a formula for the isomorphism.
Although it is often difficult to find the K-theory of an arbitrary C*-algebra, finding the K-theory of $M_{n}$ is a easy. In this case, we have the isomorphism $K_{0}(M_{n})\rightarrow\mathbb{Z}$ given by $[p]_{0}\mapsto\mathrm{tr}\hphantom{.}p$, where $\mathrm{tr}\hphantom{.}p$ is the matrix trace of $p\in\mathcal{P}_{\infty}(M_{n})$. Similarly, for a locally compact, Hausdorff, second countable, totally disconnected space $X$, the K-theory of $C_{0}(X,M_{n})$ is also described in terms of a trace operator.
To find the K-theory of the C*-algebra of an AF groupoid, we first consider the building blocks of the groupoid. In other words, we consider an elementary groupoid of the form $R(\sigma)$. We show that $C^{*}(R(\sigma))$ is locally approximated by groupoids of the form $C_{0}(V,M_{n})$. Then we apply approximation to find a formula for the $K_{0}$-group of $C^{*}(R(\sigma))$ and an isomorphism from this group to the corresponding homology group.
@@ 4. Main result
Let $p\in C_{0}(X,M_{n})$ be a projection.Let $K\subset X_{\varphi}$ be a compact subset such that $\|p(x)\|<1$ for every $x\notin K$. It follows from [34, Propositions 2.2.4 and 2.2.7] that $\mathrm{tr}\hphantom{.}p(x)=0$ for all $x\notin K$. Then $\mathrm{tr}\hphantom{.}p\in C_{c}(X_{\varphi},\mathbb{N})$ and $p\in C_{c}(X,\mathbb{N})$.
@@
Then $\widetilde{p}$ and $\widetilde{q}$ are projections in $C_{0}(X,M_{n})$. Also, $\widetilde{p}=v^{*}\widetilde{q}v$ and $v^{*}v=vv^{*}=\mathrm{id}1_{K}$. Considering the unital C*-algebra $C(K,M_{n})\subset C_{0}(X,M_{n})$, it follows from [34, Propositions 2.2.4 and 2.2.7] that there exists $u_{p},u_{q}\in C_{0}(X,M_{n})$ with support in $K$ such that
@@
    where we apply Renualt’s $j$ map [30, Proposition II.4.2] to identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$.
@@
    By [40, Proposition 9.3.1], for all $x\in X$, the $\ast$-representation $\pi_{x}:C_{c}(R(\sigma))\rightarrow B(\ell^{2}(R(\sigma)_{x}))$ is given by
@@
    Given $x\in X$, let $u=\varphi(\sigma(x))\in X_{\varphi}$. [40, Proposition 9.3.1] implies that there exists a unitary $U_{(x,u)}:\ell^{2}(R(\sigma)_{u})\rightarrow\ell^{2}(R(\sigma)_{x})$ such that $\pi_{x}=U_{(x,u)}\pi_{u}U_{(x,u)}^{*}$. Then $\|\pi_{x}(f)\|=\|\pi_{u}(f)\|$. Therefore,
@@
The following theorem gives an isomorphism of the ordered groups $K_{0}(C^{*}(G))$ and $H_{0}(G)$. Corollary 5.2 by Farsi, Kumjian, Pask and Sims [12] and Matui’s [20, Theorem 4.10] prove isomorphisms of homology groups of groupoids and the corresponding $K_{0}$-group. However, [12, Corollary 5.2] does not prove whether such isomorphism preseves positive elements from one group to another, and Matui only considers groupoids with compact unit spaces, without providing a detailed proof. Our theorem generalises both results, and we describe the isomorphism by an explicit formula.
@@
    the isomorphism $\nu$ is precisely the map given by [12, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
@@ 3. Part 3
5. Application to Deaconu-Renault groupoids
Here, we apply the ordered group isomorphism from Theorem 1 to characterise the AF embeddability of the C*-algebra of a Deaconu-Renault groupoid. First, we define this class of groupoid in Subsection 5.1, then we prove the commutative diagram 1.1 in Section 5.3, which we used in the proof of Theorem 2 in Subsection 5.4.
5.1. Deaconu-Renault groupoids
Here we use the notation from [42]. It follows from [9, Theorem 1] that the sets $\mathcal{U}_{A,B}^{m,n}$ generate a topology that makes $\mathcal{G}$ a topological groupoids which is locally compact, Hausdorff, second countable, étale. We denote the groupoid $\mathcal{G}$ endowed with this topology a Deaconu-Renault groupoid.
The first example of this groupoid was described in Renault’s book [30, Section III.2] by a class of groupoids whose C*-algebras are Cuntz algebras, where $X$ is the set of sequences $(x_{i})_{i\in\mathbb{N}}$, with $x_{i}\in\{0,1,\dots,n\}$ (or $x_{i}\in\mathbb{N}$ for the Cuntz algebra $\mathcal{O}_{\infty}$), and $\sigma$ was given by $\sigma(x_{0}x_{1}x_{2}\dots)=x_{1}x_{2}\dots$ Deaconu [9] generalises Renault’s definition by considering compact Hausdorff spaces $X$ and assumes that $\sigma:X\rightarrow X$ is a covering map. Under these assumptions, he proved that when $\sigma$ is a homeomorphism, then $C^{*}(\mathcal{G})$ is isomorphic to the crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$. Anantharaman-Delaroche [1, Example 1.2 (c)] generalised it to the case where $\sigma$ is a surjective local homeomorphism, and $X$ is locally compact. She also gave a sufficient conditions to make the C*-algebra of this groupoid purely infinite [1, Propositions 4.2 and 4.3]. Then Renault [31] generalised this definition by considering a map $\sigma$ that is defined only on an open subset of $X$. For the purpose of our paper, here we consider the definition by Anantharaman-Delaroche, and we impose the additional conditions on the space $X$ to be Hausdorff and second countable.
By [31, Proposition 2.9], the groupoid is also amenable. Note that there are different notions of amenability for groupoids: topological amenable and measurewise amenable. By Theorem 3.3.7 of [2], these two notions coincide for groupoids that are locally compact, Hausdorff and étale. In this case, [2, Proposition 6.1.8] implies that the reduced and full C*-algebras of the groupoid coincide.
Remark 5.1.
We use the notation $c^{-1}(0)$ for the groupoid $\bigcup_{n=1}^{\infty}R(\sigma^{n})$ in Example 2.5 because of the following. If $\mathcal{G}$ is the Deaconu-Renault groupoid and $c:\mathcal{G}\rightarrow\mathbb{Z}$ is given by $c(x,k,y)=k$ (in this case, $c$ is called the canonical cocycle), then $c^{-1}(0)$ is isomorphic to the subgroupoid $\{(x,0,y)\in\mathcal{G}\}$ of $\mathcal{G}$. The isomorphism is defined by $(x,y)\mapsto(x,0,y)$.
5.2. Skew-products and crossed products
Definition 5.2.
[30, Definition I.1.6] Let $G$ be a groupoid and $c:G\rightarrow\mathbb{Z}$ a homomorphism. Define the sets
respectively. Then $G(c)$ becomes a groupoid, called a skew-product.
Lemma 5.3.
Assume that $G$ is a locally compact, Hausdorff, second countable, étale, measurewise amenable groupoid, and suppose that $c:G\rightarrow\mathbb{Z}$ is a continuous cocycle. We define a dynamical system $\alpha^{c}$ induced by $c$. For every $z\in\mathbb{T}$, $f\in C_{c}(G)$ and $g\in G$, we set $\alpha_{z}^{c}(f)(g)=z^{c(g)}f(g)$. By Proposition II.5.1 of Renault’s book [30], there exists a unique dynamical system defined by the equation above. We make an abuse of notation by denoting this dynamical system $\alpha^{c}:\mathbb{T}\rightarrow\mathrm{Aut}(C^{*}(G))$. The following proposition by Renault [30, Proposition II.5.7] gives an isomorphism between the C*-algebra of skew-product groupoids and the corresponding crossed products. For an introduction to crossed products, see Chapters 1 and 2 of Dana William’s book [43].
Proposition 5.4.
[…]
for $F\in C_{c}(G(c))$, $z\in\mathbb{T}$, $g\in G$.
Now we find a formula for the action $\beta$ on $C^{*}(G(c))$ induced by the dual action $\widehat{\alpha}^{c}$. We will apply this formula in Chapter 5.4 for Deaconu-Renault groupoids in order to understand in more detail the condition $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$, given by Brown’s theorem.
Proposition 5.5.
Proof.
[…]
$=F(g,a+n).$
By definition, we have that $\rho\rtimes u\circ\widehat{\alpha}^{c}=\beta\circ\rho\rtimes u$. Then $\rho\rtimes u$ is an equivariant isomorphism from $(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T},\widehat{\alpha}^{c},\mathbb{Z})$ onto $(C^{*}(G(c)),\beta,\mathbb{Z})$. By [43, Lemma 2.65], $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic. ∎
5.3. The Commutative Diagram
@@ 5.3.1. Part 1 of the diagram
5.3.2. Part 2 of the diagram
In the second part of the diagram, we have an isomorphism between homology groups of different groupoids. We will obtain this isomorphism by applying the techniques of homological similarity of groupoids. Until the rest of this section, we fix $\varphi:X\rightarrow X$ be a continuous section of the surjective local homeomorphism $\sigma:X\rightarrow X$ that defines $\mathcal{G}$. Let us define the map that induces the similarity of the groupoids $\mathcal{G}(c)$ and $c^{-1}(0)$.
@@ 5.3.2. Part 2 of the diagram
5.4. AF embeddability for Deaconu-Renault groupoids
Here we show Theorem 2, which gives a condition on the map $\sigma:X\rightarrow X$, characterising when the corresponding groupoid C*-algebra $C^{*}(\mathcal{G})$ is AF embeddable. We prove this result by applying the commutative diagram of Section 5.3 in combination with Theorem 1.2.
@@ 5.4. AF embeddability for Deaconu-Renault groupoids
Let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the continuous cocycle given by $c(x,k,y)=k$ for $(x,k,y)\in\mathcal{G}$. By [12, Corollary 5.2], we have that $\mathcal{G}(c)$ is an AF groupoid, and therefore the C*-algebra $C^{*}(G(c))$ is AF. For simplicity of notation, we denote this C*-algebra by $B$.
@@ 6.1. Graph algebras
Given a row-finite graph with no sources, [19, Theorem 4.2] gives an isomorphism from the graph algebra $C^{*}(E)$ to the C*-algebra $C^{*}(G_{E})$ defined by $1_{e}\mapsto 1_{Z(e,r(e))}$, where
@@
The groupoid is almost identical to the groupoid in [19, Definition 2.3], but we made some changes to make it compatible with the notation in Raeburn’s book [29]. In constrast to our definition, in [19] two consecutive edges $x_{i},x_{i+1}$ satisfy $r(x_{i})=s(x_{i+1})$. Also, their groupoid has elements of the form $(x,-k,y)$ instead of $(x,k,y)$.
@@
    But if $y\neq z_{k+1}$ and $\sigma^{n}(y)=z_{k}$, then $y$ satisfies condition $(2)$ of Lemma 6.1 for $n,p\geq 1$. So $f(y)\leq 0$ by item (b) of this proof. Hence, $f(z_{k})\leq f(z_{k+1})$. It is straightforward from the definition that $z_{p}=z_{0}$. This implies that $f(z_{0})=f(z_{1})=\dots=f(z_{p})$. By (6.1), $f(y)=0$ for all $y\neq z_{1}$ such that $\sigma^{n}(y)=z_{0}=\sigma^{n}(x)$. Since $z_{1}$ satisfies condition $(2)$ of 6.1 for 00 and $p$, we have that $x\neq z_{1}$. Therefore $f(x)=0$.
@@
So far we have proved $(i)\Leftrightarrow(v)$. By Theorem 2, we have the equivalence $(i)\Leftrightarrow(ii)\Leftrightarrow(iii)\Leftrightarrow(v)$. The implication $(iii)\Rightarrow(iv)$ is straightforward, and $(iv)\Rightarrow(v)$ follows from the same argument used in [38, Theorem 1.1]. ∎